# Improving Representation of High-frequency Components for Medical Foundation Models

### v2

@@ Abstract
Foundation models have recently attracted significant attention for their impressive generalizability across diverse downstream tasks. However, these models are demonstrated to exhibit great limitations in representing high-frequency components and fine-grained details. In many medical imaging tasks, the precise representation of such information is crucial due to the inherently intricate anatomical structures, sub-visual features, and complex boundaries involved. Consequently, the limited representation of prevalent foundation models can result in significant performance degradation or even failure in these tasks. To address these challenges, we propose a novel pretraining strategy, named Frequency-advanced Representation Autoencoder (Frepa). Through high-frequency masking and low-frequency perturbation combined with adversarial learning, Frepa encourages the encoder to effectively represent and preserve high-frequency components in the image embeddings. Additionally, we introduce an innovative histogram-equalized image masking strategy, extending the Masked Autoencoder (MAE) approach beyond ViT to other architectures such as Swin Transformer and convolutional networks. We develop Frepa across nine medical modalities and validate it on 32 downstream tasks for both 2D images and 3D volume data. Without fine-tuning, Frepa can outperform other self-supervised pretraining methods and, in some cases, even surpasses task-specific trained models. This improvement is particularly significant for tasks involving fine-grained details, such as achieving up to a +15 increase in dice score for retina vessel segmentation and a +7 increase in IoU for lung nodule detection. Further experiments quantitatively reveal that Frepa enables superior high-frequency representations and preservation in the embeddings, underscoring its potential for developing more generalized and universal medical image foundation models.
@@
The limitations in high-frequency representation are primarily due to the dominant role of global information in current training strategies. For instance, autoencoder pretraining methods like the masked autoencoder (MAE) [1] typically produce reconstructions that capture only general semantic information while lacking specific details [11]. In fact, during the reconstruction process, low-frequency, global information tends to be more substantial, causing the model to emphasize these components over high-frequency details [16], resulting in limited detailed information. A similar issue is observed in contrastive learning-based methods [17]. Techniques such as CLIP [3] rely on textual information and label-level annotations [6], which contain more global-level information than pixel-level details, leading to a neglect of detailed information capturing. Additionally, as the commonly used backbone network, the vanilla vision transformer (ViT) exacerbates this issue. Compared to convolutional operators (e.g., ConViT [18]) and multi-scale feature maps (e.g., Swin Transformer, SwinT [19]), ViT has relatively weaker inductive biases, which results in a reduced ability to capture high-frequency components [20, 21]. This is also evident from Fig. 1a. Consequently, the inefficiency in capturing high-frequency information presents significant challenges for current foundation models in medical imaging tasks.
@@
Motivated by the aforementioned observations, we introduce an innovative pretraining strategy for foundational models, termed the Frequency-advanced representation autoencoder (Frepa). Frepa incorporates a novel imaging corruption and reconstruction branch designed to improve the utilization and representation of high-frequency components. This branch, termed frequency dual-component masking, includes two key steps: 1) High-frequency domain masking: We randomly mask high-frequency domains in the frequency spectrum. During training, the image encoders use the remaining information to reconstruct the masked frequency components, encouraging the encoders to learn and leverage realistic high-frequency distributions. 2) Low-frequency perturbation: To prevent the image encoders from relying excessively on easily learned low-frequency components, we introduce a zero-mean perturbation to the low-frequency part of the spectrum, prompting the encoders to rely more heavily on the high-frequency components. Additionally, we propose a hierarchical frequency-to-spatial loss and design an adversarial training strategy to achieve a more robust image reconstruction. We also modify the MAE by incorporating histogram-equalized masking methods, extending its application beyond vanilla ViT to convolutional networks and multi-scale ViT architectures. Furthermore, we integrate the 2D pretrained encoder with iterative intra- and inter-image transformer blocks, enabling its direct application to 3D volume data without architectural modifications. Frepa is developed on 17 million images and validated on 32 downstream tasks across 9 modalities. Frepa not only demonstrates superior performance in general medical tasks but also shows more significant improvement in tasks involving fine-grained information, where other models yield almost failed results. For instance, Frepa outperforms the best other models by approximately 15 in dice similarity coefficient (DSC) for retina vessel segmentation, 12 in DSC for pulmonary artery-vein segmentation, and 7 in Intersection over Union (IoU) for pulmonary nodule detection. Additionally, we also design a quantitative experiment to directly demonstrate that Frepa enables better representation and preservation of high-frequency components in the image embedding compared to other autoencoder pretraining methods. These results showcase that Frepa outperforms in capturing high-frequency components and fine-grained details without compromising the representation of global-level semantic information. Our codes are publicly available as https://github.com/Arturia-Pendragon-Iris/Frepa.
@@ 2.2 Medical foundation model
As foundation models for natural images have achieved significant success, their applications in the medical field are gaining increasing attention. Current research on medical foundation models can be broadly classified into specialist models and generalist models. Specialist models are typically designed to address clinical tasks within a single imaging modality or specific disease type. Examples include USFM [4] for ultrasound, RETFound [24] for retinal images, and EndoFM [25] for endoscopy videos. Although specialist models can achieve higher accuracy and efficiency within their specific modalities, they often face challenges in generalizing to other modalities. In contrast, generalist models, such as MedSAM [5], MedDr [6], and LLAVA [23], offer broader applicability and generalizability across various modalities and tasks. These methods rely on large amounts of training data, computational resources, and parameters, achieving progress in generalization. However, in the process of transferring pretraining techniques from natural images, these models often don’t adequately consider the data characteristics of medical images, such as fine anatomical structures, multi-scale objects, and complex, uncertain boundaries. Consequently, they may perform poorly on medical imaging tasks that depend on fine-grained details and high-frequency information [5, 6] (Fig. 1), posing a significant barrier to their practical application in clinical settings. In this work, we focus on lightweight generalist models and introduce a novel pretraining strategy designed to enhance the ability of foundation models to represent high-frequency information. Our approach aims to improve performance on fine-grained tasks, thereby extending the practical applicability and generalizability of these models in clinical practice.
@@ 3.1.1 Frequency dual-component masking
To enhance image understanding and high-frequency representation, we introduce the frequency dual-component masking strategy. This approach consists of two steps: high-frequency masking and low-frequency perturbation. By predicting image components of the masked frequencies, the image encoder can learn the high-frequency components and their realistic distributions. Simultaneously, the low-frequency perturbation prevents the encoder from overly relying on simple low-frequency components, instead encouraging it to utilize medium-frequency and high-frequency components for image reconstruction. This pretraining strategy can be seamlessly applied to both vanilla ViT, multi-scale ViTs, and convolutional networks without requiring changes to the framework (Fig. 2,a).
High-frequency domain masking Given an image , we can obtain its frequency spectrum via the 2D discrete centered Fourier transform:
@@
Low-frequency perturbation After the aforementioned stage, the high-frequency components are severely distorted. However, most low-frequency components remain unaffected because is zero at the center patch. To further reduce the dependence of the image encoder on the low-frequency components, we introduce low-frequency perturbation. Specifically, we add intensity-attenuated, zero-mean noise to the masked frequency spectrum:
Here is the Gaussian distribution, and is the standard deviation. We set to make the root mean square error (RMSE) of perturbed images maintaining approximately the same with other image masking pretraining methods [2, 26]. We also set the center of the noise to zero to ensure that the mean value of the image remains unchanged before and after the perturbation. Following these two stages, we perform the inverse Fourier transform on the corrupted frequency spectrum to obtain the degraded image.
@@ 3.1.2 Equaled-histogram image-domain masking
Although advanced multi-scale ViTs and convolutional networks have demonstrated superior high-frequency capturing abilities compared to the vanilla ViT [20, 19], the success of MAE relies on the global property and asymmetric structure of the vanilla ViT, which enables the application of the self-attention mechanism to discrete, non-overlapping image patches [27]. In contrast, multi-scale ViTs and convolutional networks typically incorporate local window operators, such as convolutional layers, which prevent them from this same capability. It has been shown that directly setting all masked patches to zero is inefficient [28], as this approach significantly distorts the image histogram distribution (Fig. 3). Additionally, the convolutional operator can result in the ”mask pattern vanishing” issue [28], causing inconsistent masking effects and ratios across different encoding layers.
To address this problem, we propose an innovative and straightforward image masking strategy, termed the histogram-equalized masking strategy. Specifically, we first uniformly sample 30 of the visible image patches, each with a size of . Unlike MAE [1], we do not discard the masked patches; instead, we replace them with random noise that has the same histogram distribution to the original patches. Denoting the one masked patch as , such an operation can be formulated as
where is the histogram-equalized noise, and denotes the histogram distribution. Practically, such distributions can be approximated by a Gaussian distribution with the same mean and variance as the original patches. Images corrupted via this strategy can keep the histogram consistency before and after the masking (Fig. 3). Notably, the frequency dual-component masking can also preserve the histogram distribution.
@@ 3.2 Pretraining pipeline of Frepa
Following the general autoencoder pretraining approach [1, 2, 10], we utilize lightweight image decoders for the network, composed of only several convolutional layers. To better guide the image encoder in leveraging high-frequency components and prevent key high-frequency details from being diluted by low-frequency components, we employ a high-frequency augmentation strategy. Specifically, we manually extract the high-frequency and edge features from the raw image via a filter operator , and then concatenated it with the degraded image along the channel dimension. Then the pretraining process can be formulated as
@@
We also incorporate the image gradient loss to restrict the reconstruction of details and textures.
@@
denotes the high-pass exponential filter, while a higher represents less preservation of the low-frequency components. This loss function enables the optimization of high-frequency reconstruction in the image domain without directly imposing restrictions on the frequency spectrum, thereby preventing the formation of artifacts.
To further enhance the representation of high-frequency components, we propose an adversarial training strategy specifically designed for frequency dual-component masking. This strategy encourages the image embeddings to remain consistent under different masking and perturbation samplings. Specially, given two different masking and perturbation sampling, denoting as and , the corresponding degraded images are and , respectively. The objective function for the adversarial training is as follows:
where denotes the image encoders and is the Jensen–Shannon divergence. The total loss functions can thus be formulated as follows:
@@ 3.3 Extending to 3D volume decoding
The network and pretraining strategy above are designed for 2D images. However, in clinical practice, many datasets consist of 3D volumes and image sequences, such as computed tomography (CT) and magnetic resonance imaging (MRI) scans. The current image encoder architecture cannot be directly applied to these types of data. To address this challenge, inspired by the design of spatial-temporal networks [32], we have developed an image-fusion architecture to seamlessly transfer pretrained 2D encoders to 3D volume data (Fig. 2 c).
Given a volume data , where and represent the number of images and channels, respectively, we first apply the encoder to each image to obtain image embeddings with feature channels and embedding size . We then incorporate the intra-image position embedding into the embeddings as , and feed them into an inter-image transformer block to extract intra-image features, resulting in output , where is the token dimension. The transformer block follows the commonly used architecture [33]. is resized to and then processed by an inter-image transformer block, incorporating an inter-slice position embedding , to fuse the extracted features between contextual images. These two transformer blocks are iteratively applied to the embeddings to integrate intra- and inter-image information (with three iterations in our setting). The final output is then passed to a decoder head for various downstream tasks.
| Modality | Dataset | Metric | MAE | MFM | CLIP | MedSAM | MedDr | Frepa+ViT | Frepa+SwinT |
| CT | Chest cancer[34] | ACC | 0.888 | 0.794 | 0.889 | 0.857 | - | 0.937 | 0.920 |
| AUC | 0.983 | 0.964 | 0.982 | 0.973 | - | 0.991 | 0.983 | ||
| X-Ray | COVID19 Radiography[35] | ACC | 0.962 | 0.949 | 0.966 | 0.966 | - | 0.964 | 0.968 |
| AUC | 0.999 | 0.996 | 0.997 | 0.997 | - | 0.998 | 0.998 | ||
| Pneumonia MNIST[12] | ACC | 0.919 | 0.918 | 0.939 | 0.903 | 0.870 | 0.934 | 0.932 | |
| AUC | 0.973 | 0.966 | 0.982 | 0.965 | - | 0.970 | 0.981 | ||
| COVID19 PA Dataset[36] | ACC | 0.934 | 0.908 | 0.943 | 0.897 | - | 0.939 | 0.941 | |
| AUC | 0.980 | 0.980 | 0.981 | 0.977 | - | 0.984 | 0.984 | ||
| Lung cancer[37] | Detection | 0.911 | 0.908 | 0.843 | 0.899 | - | 0.951 | 0.948 | |
| IoU | 0.596 | 0.607 | 0.482 | 0.566 | - | 0.656 | 0.677 | ||
| OCT | OCT2017[38] | ACC | 0.815 | 0.866 | 0.868 | 0.837 | - | 0.888 | 0.880 |
| AUC | 0.932 | 0.960 | 0.972 | 0.944 | - | 0.970 | 0.963 | ||
| OCT MNIST[12] | ACC | 0.708 | 0.737 | 0.670 | 0.661 | 0.675 | 0.793 | 0.786 | |
| AUC | 0.903 | 0.913 | 0.908 | 0.880 | - | 0.940 | 0.932 | ||
| Ultrasound | Fetal Plane[39] | ACC | 0.928 | 0.922 | 0.932 | 0.919 | - | 0.897 | 0.919 |
| AUC | 0.959 | 0.959 | 0.946 | 0.955 | - | 0.945 | 0.951 | ||
| Breast MNIST[12] | ACC | 0.852 | 0.814 | 0.852 | 0.794 | 0.718 | 0.846 | 0.846 | |
| AUC | 0.861 | 0.834 | 0.833 | 0.823 | - | 0.863 | 0.798 | ||
| Retina | Aptos2019[40] | ACC | 0.764 | 0.796 | 0.758 | 0.788 | - | 0.802 | 0.808 |
| AUC | 0.882 | 0.908 | 0.891 | 0.893 | - | 0.914 | 0.906 | ||
| Dermatoscope | ISIC2019[41] | ACC | 0.701 | 0.696 | 0.705 | 0.701 | - | 0.715 | 0.717 |
| AUC | 0.887 | 0.902 | 0.892 | 0.891 | - | 0.921 | 0.916 |
| Modality | Dataset | Metric | MAE | MFM | CLIP | MedSAM | MedDr | Frepa+ViT | Frepa+SwinT |
| CT | COPD# | ACC | 0.869 | 0.828 | 0.878 | 0.801 | - | 0.887 | 0.882 |
| AUC | 0.946 | 0.933 | 0.930 | 0.929 | - | 0.954 | 0.960 | ||
| Pneumonia# | ACC | 0.867 | 0.842 | 0.897 | 0.882 | - | 0.897 | 0.887 | |
| AUC | 0.924 | 0.913 | 0.933 | 0.924 | - | 0.935 | 0.929 | ||
| Luna16 Lung Nodule[42] | Detection | 0.885 | 0.872 | 0.775 | 0.876 | - | 0.905 | 0.920 | |
| IoU | 0.569 | 0.548 | 0.410 | 0.518 | - | 0.626 | 0.633 |

### v3

# Improving Representation of High-frequency Components for Medical Visual Foundation Models

@@ Abstract
Foundation models have recently attracted significant attention for their impressive generalizability across diverse downstream tasks. However, these models are demonstrated to exhibit great limitations in representing high-frequency components and fine-grained details. In many medical imaging tasks, the precise representation of such information is crucial due to the inherently intricate anatomical structures, sub-visual features, and complex boundaries involved. Consequently, the limited representation of prevalent foundation models can result in significant performance degradation or even failure in these tasks. To address these challenges, we propose a novel pretraining strategy for both 2D images and 3D volumes, named Frequency-advanced Representation Autoencoder (Frepa). Through high-frequency masking and low-frequency perturbation combined with embedding consistency learning, Frepa encourages the encoder to effectively represent and preserve high-frequency components in the image embeddings. Additionally, we introduce an innovative histogram-equalized image masking strategy, extending the Masked Autoencoder (MAE) approach beyond ViT to other architectures such as Swin Transformer and convolutional networks. We develop Frepa across nine medical modalities and validate it on 32 downstream tasks for both 2D images and 3D volume data. Without fine-tuning, Frepa can outperform other self-supervised pretraining methods and, in some cases, even surpasses task-specific foundation models. This improvement is particularly significant for tasks involving fine-grained details, such as achieving up to a +15 increase in dice score for retina vessel segmentation and a +8 increase in IoU for lung tumor detection. Further experiment quantitatively reveals that Frepa enables superior high-frequency representations and preservation in the embeddings, underscoring its potential for developing more generalized and universal medical image foundation models.
@@
The limitations in high-frequency representation are primarily due to the dominant role of global information in current training strategies. For instance, autoencoder pretraining methods like the masked autoencoder (MAE) [1] typically produce reconstructions that capture only general semantic information while lacking specific details [11]. In fact, during the reconstruction process, low-frequency, global information tends to be more substantial, causing the model to emphasize these components over high-frequency details [16], resulting in limited detailed information. A similar issue is observed in contrastive learning-based methods [17]. Techniques such as CLIP [3] rely on textual information and label-level annotations [6], which contain more global-level information than pixel-level details, leading to a neglect of detailed information capturing. Additionally, as the commonly used backbone network, the vanilla vision transformer (ViT) exacerbates this issue. Compared to convolutional operators (e.g., ConvNeXt [18]) and multi-scale feature maps (e.g., Swin Transformer, SwinT [19]), ViT has relatively weaker inductive biases, which results in a reduced ability to capture high-frequency components [20, 21]. This is also evident from Fig. 1a. Consequently, the inefficiency in capturing high-frequency information presents significant challenges for current foundation models in medical imaging tasks.
@@
Motivated by the aforementioned observations, we introduce an innovative pretraining strategy for foundational models, which is applicable for both 2D images and 3D volumes, termed the Frequency-advanced representation autoencoder (Frepa). Frepa incorporates a novel imaging corruption and reconstruction branch designed to improve the utilization and representation of high-frequency components. This branch, termed frequency dual-component masking, includes two key steps: 1) High-frequency domain masking: We randomly mask high-frequency domains in the frequency spectrum. During training, the image encoders use the remaining information to reconstruct the masked frequency components, encouraging the encoders to learn and leverage realistic high-frequency distributions. 2) Low-frequency perturbation: To prevent the image encoders from relying excessively on easily learned low-frequency components, we introduce a zero-mean perturbation to the low-frequency part of the spectrum, prompting the encoders to rely more heavily on the high-frequency components. Additionally, we propose a hierarchical frequency-to-spatial loss and design an embedding learning strategy to achieve a more robust image reconstruction. We also modify the MAE by incorporating histogram-equalized masking methods, extending its application beyond vanilla ViT to convolutional networks and multi-scale ViT architectures. Furthermore, we propose to integrate the 2D pretrained encoder with iterative intra- and inter-image transformer blocks, enabling it direct application to 3D volume data without retraining. Frepa is developed on 17 million images and validated on 32 downstream tasks across 9 modalities. Frepa not only demonstrates superior performance in general medical tasks but also shows more significant improvement in tasks involving fine-grained information, where other models yield almost failed results. For instance, Frepa outperforms the best other models by approximately 15 in dice similarity coefficient (DSC) for retina vessel segmentation, 12 in DSC for pulmonary artery-vein segmentation, and 8 in Intersection over Union (IoU) for lung tumor detection. Additionally, we also design a quantitative experiment to directly demonstrate that Frepa enables better representation and preservation of high-frequency components in the image embedding compared to other autoencoder pretraining methods. These results showcase that Frepa outperforms in capturing high-frequency components and fine-grained details without compromising the representation of global-level semantic information. Our codes are publicly available as https://github.com/Arturia-Pendragon-Iris/Frepa.
@@ 2.2 Medical foundation model
As foundation models for natural images have achieved significant success, their applications in the medical field are gaining increasing attention. Current research on medical foundation models can be broadly classified into specialist models and generalist models. Specialist models are typically designed to address clinical tasks within a single imaging modality or specific disease type. Examples include USFM [4] for ultrasound, RETFound [24] for retinal images, SkinGPT-4 [25] for skin disease, M3FM [26] and CT-CLIP [27] for chest computed tomography, and EndoFM [28] for endoscopy videos. These models are typically designed upon the unique characteristics of their respective modalities and are capable of handling diverse downstream tasks. By leveraging knowledge learned during the pretraining phase on large-scale datasets, specialist models often outperform single-modal task-specific models. They can also be flexibly adapted to new tasks using only a small dataset [26], making them highly efficient for medical applications.
While specialist models often achieve higher accuracy and efficiency within specific modalities, they frequently struggle to generalize across different modalities. In contrast, generalist models demonstrate better generalizability across diverse modalities. Early studies have explored the application of general foundation models, such as GPT-4 [29], for clinical findings and diagnosis. Subsequently, medical generalist models designed for computer-aided diagnosis, such as BiomedGPT [30], MedDr [6], Medical Gemini [31], and LLAVA [23], were developed using public medical datasets to scale and integrate clinical knowledge. These models have shown robust capabilities in disease diagnosis, question answering, and medical report generation. Such approaches typically rely on large-scale training data, extensive computational resources, and numerous parameters, driving advancements in generalization. Additionally, generalist models targeting other tasks have also achieved notable success, including MedSAM [5] for semantic segmentation and MINIM [32] for image generation. However, in the process of transferring pretraining techniques, these models often do not adequately consider the data characteristics of medical images, such as fine anatomical structures, multi-scale objects, and complex, uncertain boundaries. Consequently, they may perform poorly on medical imaging tasks that depend on fine-grained details and high-frequency information [5, 6] (Fig. 1), posing a significant barrier to their practical application in clinical settings. In this work, we focus on lightweight generalist models and introduce a novel pretraining strategy designed to enhance the ability of foundation models to represent high-frequency information. Our approach aims to improve performance on fine-grained tasks, thereby extending the practical applicability and generalizability of these models in clinical practice.
@@ 3.1.1 Frequency dual-component masking
To enhance image representation of high-frequency information, we introduce the frequency dual-component masking strategy. This approach consists of two steps: high-frequency masking and low-frequency perturbation. By predicting image components of the masked frequencies, the image encoder can learn the high-frequency components and their distributions. Simultaneously, low-frequency perturbation prevents the encoder from overly relying on simple low-frequency components, instead encouraging it to utilize medium-frequency and high-frequency components for image reconstruction. This pretraining strategy can be seamlessly applied to both vanilla ViT, multi-scale ViTs, and convolutional networks without requiring changes to the framework (Fig. 2, a).
High-frequency domain masking Given a 2D image , we can obtain its frequency spectrum via the 2D discrete centered Fourier transform:
@@
Low-frequency perturbation After the aforementioned stage, the high-frequency components are severely distorted. However, most low-frequency components remain unaffected, as is much lower in the low-frequency domain. To prevent image restoration from relying solely on low-frequency components and instead promote understanding of high-frequency components, we introduce the low-frequency perturbation strategy. Specifically, we add intensity-attenuated, zero-mean noise to the masked frequency spectrum:
Here is the Gaussian distribution, and is the standard deviation. We set to make the root mean square error (RMSE) of perturbed images maintaining approximately the same with other image masking pretraining methods ([2, 33]). We also set the center of the noise to zero to ensure that the mean value of the image remains unchanged before and after the perturbation. Following these two stages, we perform the inverse Fourier transform on the corrupted frequency spectrum to obtain the degraded image.
@@ 3.1.2 Equaled-histogram image-domain masking
Although advanced multi-scale ViTs and convolutional networks have demonstrated superior high-frequency capturing abilities compared to the vanilla ViT [20, 19], the success of MAE relies on the global property and asymmetric structure of the vanilla ViT, which enables the application of the self-attention mechanism to discrete, non-overlapping image patches [34]. In contrast, multi-scale ViTs and convolutional networks typically incorporate local window operators, such as convolutional layers, which prevent them from this same capability. It has been shown that directly setting all masked patches to zero is inefficient [35], as this approach significantly distorts the image histogram distribution (Fig. 3). Additionally, the convolutional operator can result in the “mask pattern vanishing” issue [35], causing inconsistent masking effects and ratios across different encoding layers.
To address this problem, we propose an innovative and straightforward image masking strategy, termed the histogram-equalized masking strategy. Specifically, we first uniformly sample 70 of the visible image patches, each with a size of . Unlike MAE [1], we do not discard the masked patches; instead, we replace them with random noise that has the same histogram distribution to the original patches. Denoting the one masked patch as , such an operation can be formulated as
where is the histogram-equalized noise, and denotes the histogram distribution. Practically, such distributions can be approximated by a Gaussian distribution with the same mean and variance as the original patches. Images corrupted via this strategy can keep the histogram consistency before and after the masking, as shown in Fig. 3. Notably, this figure also shows that the frequency dual-component masking can also preserve the histogram distribution.
@@ 3.2 Pretraining pipeline of Frepa
We employ lightweight image decoders for the network, consisting of only a few convolutional layers. Compared to the Transformer decoder used in MAE [1], convolutional decoders offer significantly lower computational complexity [36], particularly given that medical images often have relatively large dimensions (e.g., in our experiments). Additionally, convolutional decoders demonstrate better capability in capturing local fine-grained details [20], aligning well with the objectives of our reconstruction task. To better guide the image encoder in leveraging high-frequency components and prevent key high-frequency details from being diluted by low-frequency components, we employ a high-frequency augmentation strategy. Specifically, we manually extract the high-frequency and edge features from the raw image via a filter operator , and then concatenated it with the degraded image along the channel dimension. Then the pretraining process can be formulated as
@@
However, this constraint may be insufficient to ensure similarity in image details and high-frequency components, as RMSE primarily focuses on overall pixel-wise intensity differences. Therefore, we additionally incorporate two loss functions, image gradient loss, and frequency-to-spatial loss, to better guide the reconstruction of fine-grained details and high-frequency components. The image gradient loss restricts the reconstruction of structures and details [38].
@@
denotes the high-pass exponential filter, while a higher represents less preservation of the low-frequency components. In our implementation, the cut-off distance of the exponential filter is set to , with ranging from 1 to 5. This hierarchical decomposition enables the optimization of high-frequency reconstruction in the image domain without directly imposing restrictions on the frequency spectrum, thereby preventing the formation of artifacts.
We also incorporate a training strategy to enhance the representation consistency, which encourages the image embeddings to remain consistent under different masking and perturbation samplings. Specifically, for a given input image , we generate two degraded versions by applying distinct masking and perturbation configurations, and , and the corresponding degraded images are denoted as and , respectively. The consistency objective aims to make the embeddings of these two degraded images as similar as possible, and we use the Jensen-Shannon (JS) divergence to evaluate their distribution divergence:
where denotes the image encoder. The total loss functions can thus be formulated as follows:
@@ 3.3 Extending to 3D volume
In clinical practice, many datasets are composed of 3D volumes and image sequences, such as computed tomography (CT) and magnetic resonance imaging (MRI) scans. The pretraining strategy can be readily adapted for 3D volume data or applied by transferring pretrained 2D encoders directly to 3D volumes (Fig. 2c).
3.3.1 Adapt pretraining strategy for 3D volume
Given a 3D volume consisting of stacked 2D images, similar masking strategies can be applied. For frequency dual-component masking, we first compute its 3D frequency spectrum and apply high-frequency domain masking by modifying Eq. 2 into a 3D manner
Here is the masking probability of 3D patch , while denotes the 3D Euclidean distance from the center of patch to the frequency spectrum center. Low-frequency perturbation can also be implemented by adjusting the Gaussian distribution in Eq.3 to 3D distribution. In the equaled-histogram image-domain masking, we set the patch size to [41] and perform masking accordingly,
Here is the masked patch in the 3D volume. The pretraining pipeline remains the same with the 2D operations.
3.3.2 Transfer pretrained 2D encoder to 3D volumes
Although Frepa can be directly adapted for 3D volume data, re-pretraining can cost extensive computational resources. Here we propose a method that can transfer pretrained 2D encoders directly to 3D volumes without re-training. Specifically, given a volume data , we first apply the pretrained 2D encoder to each stacked image to obtain embeddings with feature channels and embedding size . To preserve the spatial information within each image, fixed sinusoidal positional embedding is added to the rearranged embeddings as . The resulting embeddings are then processed by an intra-image spatial transformer block, which employs a multi-head self-attention mechanism and a feedforward network (following the standard transformer architecture [42]) to capture spatial relationships within each individual image. This step results in output embeddings , where denotes the token dimension.
The embeddings are subsequently reshaped to , enabling the transformer to focus on relationships between corresponding patches across the stacked images. To encode the relative positions of slices within the stack, an inter-image positional embedding is added to the rearranged embeddings. The resulting representation is then processed by an inter-image transformer block (same architecture as the intra-image transformer block), which captures contextual relationships between the stacked images by attending to corresponding patches across slices. The intra- and inter-image transformer blocks are applied iteratively (three iterations in the current implementation) to progressively integrate intra-image and inter-image information, thereby refining the representation of both local and global patterns. Finally, the output embeddings are passed to a decoder head tailored for various downstream tasks. During the training phase, the parameters of the pretrained 2D encoder are frozen, while all parameters within the transformer blocks are optimized.